OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 3x + y = 14 -3x + 2y = 1 5 3 y can be any number as there are infinitely many solutions to the system. There is no y value as there is no solution to the system.

OpenStudy (anonymous):

MEDAL!!!!

OpenStudy (anonymous):

http://www.purplemath.com/modules/systlin5.htm

OpenStudy (anonymous):

I don't see my problem on there what's the answer

OpenStudy (anonymous):

I put you an example. So YOU could solve your problem. And if you get stuck I could help you : )

OpenStudy (anonymous):

Ok

OpenStudy (mayankdevnani):

I can tell you how to solve two equation with two variables

OpenStudy (anonymous):

I got 3 for my answer

OpenStudy (mayankdevnani):

\[\large \bf 3x+y=14................(1)\] \[\large \bf -3x+2y=1..............(2)\] From equation (1), \[\large \bf y=14-3x.........................(3)\] now put equation (3) in equation (2), \[\large \bf -3x+2(14-3x)=1\] \[\large \bf -3x+28-6x=1\] \[\large \bf -9x=-27\] so, \[\large \bf x=3\] Now,put the value of x=3 in the equation (3) and then find the value of y, \[\large \bf y=14-3(3)=14-9=5\] So,we get \[\large \bf x=3~and~y=5~as~\color{red}{Answers}\]

OpenStudy (mayankdevnani):

understood? @kaylamarie081

OpenStudy (anonymous):

Yes very much . thank you !

OpenStudy (mayankdevnani):

and you also @charliealbert . I hope you both understand very well !!

OpenStudy (anonymous):

Yes, sure danke

OpenStudy (anonymous):

So answer is 5

OpenStudy (mayankdevnani):

yeah